# Robust Variational Physics-Informed Neural Networks

### 11 Dec 2023 v2

@@ 5 Numerical examples
To illustrate the proposed strategy’s performance, we consider several diffusion-advection problems subject to Dirichlet-type boundary conditions. We focus on problems that either are challenging for the standard finite element method (advection-dominated problems) or do not admit a solution belonging to $H^{2}(\Omega)$. To construct the loss functionals, we consider two distinct discrete test spaces: one employing standard FE piece-wise polynomials and another employing spectral orthonormal test functions. Here, we do not intend to compare the performance between different test spaces, but rather our aim is to numerically validate the robustness of RVPINNs.
@@ 5.1 Diffusion-advection model problem
$a(u\,,v):=\left(\mathbf{\varepsilon}\nabla u-\mathbf{\beta}\,u\,,\nabla v\right)_{0},\quad\text{and}\quad l(v)=\langle f\,,v\rangle,$ (40)
@@ 5.2 Discrete setting
For simplicity, we consider only one-dimensional problems. We set $\Omega:=(-1,1)$ and define two kinds of test spaces. First, we consider $V_{M}$ as FE space of the standard globally continuous and piece-wise linear functions that belong to $H_{0}^{1}(\Omega)$, defined over a uniform mesh partition of $\Omega$ into $M+1$ elements $[x_{i},x_{i+1}]$, with $x_{0}=-1<x_{1}<\dots<x_{M}<x_{M+1}=1$. The corresponding Gram matrix is constructed using the inner product, inducing the norm $\|\cdot\|_{V}$, defined in (41). The second test space is intended to show examples involving an orthonormal basis for the discrete space $V_{M}$. Here, we normalize with respect to the inner product (41) the first $M$ eigenfunctions of the Laplacian satisfying the Dirichlet homogeneous boundary conditions. The $m$-th eigenfunction and the corresponding orthonormal set, respectively, are given by
@@
Therefore, the loss functional (26) is explicitly given, up to the constraint $C(u_{\theta})$, as:
@@
$\mathcal{L}^{\phi}_{r}(u_{\theta})=\dfrac{1}{\varepsilon\pi^{2}}\sum_{n=1}^{N}\sum_{m=1}^{M}\dfrac{1}{n^{2}+m^{2}}\,r\left(u_{\theta}\,,s_{m,n}\right)^{2}+\,C(u_{\theta}).$ (49)
For the architecture of the DNN approximator $u_{\theta}$, we adopt the following structure for the 1D problems: a fixed feed-forward fully connected network comprising five layers, each with 25 neurons. Throughout all layers, the activation function employed is $\tanh$. When strongly enforcing the homogeneous Dirichlet boundary conditions, we multiply the last layer’s output by $(x+1)(x-1)$. Another version includes the constrained imposition of homogeneous Dirichlet boundary conditions by selecting:
$C(u_{\theta})=|u_{\theta}(-1)|^{2}+|u_{\theta}(1)|^{2}.$ (50)
As our choice for the nonlinear solver, we opt for the ADAM optimizer (refer to [32]) and initialize it with a learning rate of $0.0005$. Our nonlinear solver undergoes up to $6000$ iterations (epochs). To approximate the integral terms appearing in the loss functional definitions of the 1D problems, we employ (a) a fixed Gaussian quadrature rule with five nodes per element when considering FE functions, and (b) a trapezoidal rule with $4000$ equally-distributed nodes when considering spectral functions. Lastly, the error $|u-u_{\theta}|_{U}$ is numerically estimated using a trapezoidal rule, with $10000$ equally-distributed nodes. For the 2D example, we consider a fixed feed-forward fully connected network comprising four layers, each with 40 neurons and the same type of activation function as in the 1D cases. For the training, $500\times 500$ equally-distributed integration points, while equally-distributed $1000\times 1000$ integration points for estimating the error $|u-u_{\theta}|_{U}$. We strongly impose the Dirichlet boundary conditions by multiplying the last layer’s output by $x(x-1)y(y-1)$.
@@ 5.3 Smooth diffusion problems
Figure 1 shows the numerical results obtained with RVPINNs with $50$ sinusoidal test functions and strong imposition of the BCs. We observe a good performance of the RVPINNs strategy, obtaining an accurate approximation of the analytical solution, as shown in Figure 0(a). Moreover, we observe a perfect match between the quantities $\|u-u_{\theta}\|_{U}$, $\|\phi\|_{V}$, and $\sqrt{\mathcal{L}^{\phi}_{r}(u_{\theta})}$ as shown in figures 0(b) and  0(c) where, in particular, the last one shows a perfect correlation between the true error and the square root of the loss functional. Figure 2 shows an almost identical behavior when considering $100$ FE test functions and strong imposition of BCs. We obtain similar results when considering the constrained imposition of BCs, but we omit them here for brevity.
Note that in the case of pure diffusive problems, the continuity constant $\mu$ and the inf-sup stability constant $\alpha$ are equal to one when considering the $H^{1}$-seminorm (cf. (42) and (45)). Therefore, $\|\phi\|_{V}$ always defines a lower bound for the error (cf. (34)) and an upper bound up to a constant close to one if the oscillation term in (35) is sufficiently small. These theoretical considerations are confirmed in figures 1 and 2
@@
To conclude this section, in Figure 4 we show the results obtained for a 2D pure diffusion example, set in the domain $\Omega=(0,1)^{2}$ analogously so the analytical solution coincides with
@@
For the training, we consider the spectral loss functional (48), with $M=N=30$ and strong imposition of the BCs. We observe a similar behavior, up to some oscillation due to the inexact integration, to the 1D case, showing in particular that the strategy is dimension-independent, as expected.
@@ 5.4 Delta source problem
which is only a $C^{0}(\Omega)$ solution, while the activation function $\tanh$ is smooth. Therefore, a major dominance of the oscillation term is expected, even when considering a sufficiently large number of test functions, as appreciated in Figure 5 where we have considered $100$ FE test functions and strong imposition of the BCs. Figure 6 shows similar results for the same configuration but with the constrained imposition of the BCs. In both cases, a good representation of the analytical solution is obtained despite being approximated with a smooth function. Additionally, figures 5 and  6 show that the loss functional defines a robust estimation of the error in the pre-asymptotic regime (before the dominance of the oscillation term). A similar behavior is obtained when considering the loss functional with spectral test functions– omitted here for brevity.
@@
Figure 6: RVPINNs approximation of the delta diffusion problem; constrained BCs imposition, $100$ FE test functions.
@@ 5.5 Advection-dominated diffusion problem
exhibiting a strong gradient near the right boundary when $\varepsilon\rightarrow 0$ (cf. figures 6(a) and 8(a)). We restrict the discussion to spectral test functions, as we observe similar behaviors when considering FE test functions.
Figure 7 shows the results obtained for $\varepsilon=0.1$, considering $50$ test functions and strong imposition of the BCs. Contrary to the pure diffusive case, Figure 6(b) reflects that the lines representing the error and the square root of the loss functional (respectively, the norm of the discrete Riesz representative) do not overlap. This behavior is expected as, in this case, the continuity constant $\mu$ is not equal to one. It deteriorates when $\varepsilon\rightarrow 0$ (cf. (42)). We also observe a decay in the expected optimal correlation between the error and the square root of the loss starting near iteration $3000$, apparently due to stability issues. This suboptimality is improved when considering the constrained imposition of the boundary conditions, as shown in Figure 8. This constrained imposition of BCs could thus serve as a stabilization technique. This is further confirmed in Figure 9, where we show the results obtained for $\varepsilon=0.005$, with $200$ test basis functions and constrained BCs. When considering instead the strong imposition of the BCs and the same number of test functions, we do not observe convergence within the first $6000$ iterations–omitted here for brevity. We observe that when considering the constrained BCs, the loss functional exhibits some differences with the norm of the Riesz representative. However, this also agrees with our findings, as the constraint does not depend on the test space. Thus, it is not represented by the Riesz representative, as the boundary is not taken into consideration in the norm definition of the test space. Nevertheless, in this case, we still observe a good correlation between the loss and the true error.
@@
Figure 8: RVPINNs approximation of the diffusion-advection problem with $\varepsilon=0.1$; constrained BCs imposition, $50$ spectral test functions.
@@
Figure 9: RVPINNs approximation of the diffusion-advection problem with $\varepsilon=0.005$; constrained BCs imposition, $200$ spectral test functions.

### 05 Mar 2024 v3

@@ 4.2 Energy norm error estimates based on local semi-discrete inf-sup condition
4.3 Error estimates in the quasi-minimizer sense
In the previous section, we derived a priori and a posteriori error estimates near the convergence regime by assuming the existence of a local solution of the Petrov-Galerkin-type problem (11). However, as mentioned in Remark (2), this assumption cannot be satisfied in several practical scenarios. Additionally, we notice that our a priori error estimate does not take into consideration the contribution of the constraint $C(\cdot)$, since it is not necessarily true that a solution $u_{\theta^{\ast}}$ of the corresponding Petrov-Galerkin-type problem will also satisfy that $C(u_{\theta^{\ast}})$ will be bounded (up to a constant) by the infimum of $C(u_{\theta})$. This is the case when $C(\cdot)$ is defined in terms of noisy data and the DNN architecture allows for functions $u_{\theta}$ satisfying $C(u_{\theta})=0$. Thus, a more general framework allowing for a less restrictive assumption must be considered. To this end, following [10], we start by considering the following definition:
Definition 1 (Quasi-minimizers).
Let $\mathcal{L}:U\mapsto\mathbb{R}$ be a cost functional. Let $\delta_{S}>0$ and $U_{NN}\subset U$ be the manifold introduced in (10), represented by a set $\theta\in\mathbb{R}^{S}$ of trainable parameters. A function $u_{\theta^{S}}\in U_{NN}$ is said to be a quasi-minimizer of $\mathcal{L}$ if the following holds true:
$\mathcal{L}(u_{\theta^{S}})\leq\inf_{\theta\in\mathbb{R}^{S}}\mathcal{L}(u_{\theta})+\delta_{S}.$ (38)
The following is a direct consequence of the RVPINN cost function being positive-valued (cf. Eq. (24)).
Proposition 7.
For a given DNN structure $U_{NN}$, there exists a sufficiently small $\delta^{\ast}\geq 0$ and a function $u_{\theta^{\ast}}\in U_{NN}$ such that, for all $\delta_{S}>\delta^{\ast}$, $u_{\theta^{\ast}}$ is a quasi-minimizer of the cost functional (18).
The following Lemma is a direct consequence of the quasi-minimizer definition and the property of the Riesz representative (17) (cf. proof of Theorem 2):
Lemma 8.
Let $u_{\theta^{\ast}}$ be a quasi-minimizer and $\delta^{\ast}>0$ be the smallest constant satisfying (38). It holds
$\left(\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-u_{\theta^{\ast}},v_{M})}{\|v_{M}\|_{V}}\right)^{2}+C(u_{\theta^{\ast}})\leq\left(\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-u_{\theta},v_{M})}{\|v_{M}\|_{V}}\right)^{2}+C(u_{\theta})+\delta^{\ast},\quad\forall\,u_{\theta}\in U_{NN}.$ (39)
We can now derive an a priori error estimate in the quasi-minimizer sense by considering the following relaxation of the Fortin’s Assumption 1444This is indeed a relaxation of Assumption 1, since in the case where there exists a solution of the Petrov-Galerkin-type problem (11) and $C(u_{\theta})=0$, we can consider $\delta_{S}=0$ and $\theta_{S}=\theta^{\ast}$..
Assumption 2 (Local Fortin’s condition II).
Given $\delta_{S}\geq 0$ such that $u_{\theta_{S}}\in U_{NN}$ is a quasi-minimizer of (18), there exists $R>0$ such that, for all $\theta\in B(\theta_{S},R)$, an operator $\Pi_{\theta}:V\mapsto V_{M}$, and a $\theta$-independent constant $C_{\Pi}>0$, verifying the same properties of Assumption 1.
Proposition 9 (Local a priori error estimate II).
Let $u_{\theta^{\ast}}$ be a quasi-minimizer and $\delta^{\ast}>0$ be the smallest constant satisfying (38). If Assumption 2 is satisfied, it holds:
$\|u-u_{\theta^{\ast}}\|^{2}_{U}+\left(\dfrac{\sqrt{2}C_{\Pi}}{\alpha}\right)^{2}C(u_{\theta^{\ast}})\leq\inf_{u_{\theta}\in U_{NN}^{\theta^{\ast}}}\left\{2\left(1+\left(\dfrac{2\mu C_{\Pi}}{\alpha}\right)^{2}\right)\|u-u_{\theta}\|^{2}_{U}+\left(\dfrac{\sqrt{2}C_{\Pi}}{\alpha}\right)^{2}\big{(}C(u_{\theta})+\delta^{\ast}\big{)}\right\}.$ (40)
Proof.
The proof is similar as in Proposition 6 by noticing that, for all $u_{\theta}\in U_{NN}^{\theta^{\ast}}$, with $U_{NN}^{\theta^{\ast}}$ defined as in (36), it holds
$\left(\dfrac{\alpha}{2C_{\Pi}}\right)^{2}\|u_{\theta^{\ast}}-u_{\theta}\|^{2}_{U}+\dfrac{1}{2}C(u_{\theta^{\ast}})&\displaystyle\leq\frac{1}{4}\left(\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u_{\theta^{\ast}}-u_{\theta},v_{M})}{\|v_{M}\|_{V}}\right)^{2}+\dfrac{1}{2}C(u_{\theta^{\ast}})\\ &\displaystyle\leq\frac{1}{4}\left(\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-u_{\theta},v_{M})}{\|v_{M}\|_{V}}+\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-u_{\theta^{\ast}},v_{M})}{\|v_{M}\|_{V}}\right)^{2}+\dfrac{1}{2}C(u_{\theta^{\ast}})\\ &\displaystyle\leq\left(\sup_{0\neq v_{M}\in V_{M}}\dfrac{a(u-u_{\theta},v_{M})}{\|v_{M}\|_{V}}\right)^{2}+\dfrac{1}{2}\left(C(u_{\theta})+\delta^{\ast}\right),\end{array}$
where the last inequality follows from the Young’s inequality $2xy\leq x^{2}+y^{2},\,\forall x,y\in\mathbb{R}$, and Lemma 8. ∎
@@ 5 Numerical examples
To illustrate the proposed strategy’s performance, we consider several diffusion-advection problems subject to Dirichlet-type boundary conditions. We focus on problems that either are challenging for the standard finite element method (advection-dominated problems) or do not admit a solution belonging to $H^{2}(\Omega)$. We consider two distinct discrete test spaces to construct the loss functionals: one employing standard FE piece-wise polynomials and another employing spectral orthonormal test functions. Here, we do not intend to compare the performance between different test spaces, but rather, we aim to numerically validate the robustness of RVPINNs.
@@ 5.1 Diffusion-advection model problem
$a(u\,,v):={\left(\mathbf{\varepsilon}\nabla u\,,\nabla v\right)_{0}+\left(\mathbf{\beta}\cdot\nabla u\,,v\right)_{0}},\quad\text{and}\quad l(v)=\langle f\,,v\rangle,$ (43)
@@
5.2 A weak BCs imposition for 1D problems
Setting $\Omega=(-1,1)$, $U=H^{1}(\Omega)$, $V=H_{0}^{1}(\Omega)$, and $\widehat{u}=(u(-1),u(1))$, the variational formulation is defined as follows [^6] Notice that this way of imposing BCs can also be applied to other alternative variational formulations. For instance, the Strong VF in the spirit of Remark 11.
$\text{Find }u\in U:a(u,(v,\widehat{v})):=(\varepsilon u^{\prime}\,,v^{\prime})_{0}+(\beta u^{\prime}\,,v)_{0}+\widehat{u}\cdot\widehat{v}=l((v,\widehat{v})):=\langle f\,,v\rangle,\,\forall\,(v,\widehat{v})\in V\times\mathbb{R}^{2}.$ (49)
Variational formulation  (49) is well-possed. Indeed, defining $\|\cdot\|^{2}_{V}:=\varepsilon\,(\cdot^{\prime}\,,\cdot^{\prime})_{0}$ and denoting by $\|\cdot\|_{E}$ the Euclidean norm of $\mathbb{R}^{2}$, we can consider the norms
$\|u\|_{U}^{2}:=\|u\|_{V}^{2}+\|\widehat{u}\|_{E}^{2}\quad\text{and}\quad\|(v,\widehat{v})\|_{V\times\mathbb{R}^{2}}^{2}:=\|v\|_{V}^{2}+\|\widehat{v}\|_{E}^{2},$ (50)
for the trial and test space, respectively. A straightforward verification proves that the boundedness condition (2) is satisfied with the upper bound constant $\mu$ of Eq. (45). It is also immediate to verify that the condition (4) is satisfied. Indeed, noticing that $(\beta v^{\prime},v)_{0}=0$, for all $v\in V$ (cf. Eq. (46)), if $a(u,(v,\widehat{v}))=0$ holds for all $u\in U$, then it holds for $u=v+u_{\widehat{v}}$, with the last being the (unique) solution of the problem:
Find $\text{Find }u_{\widehat{v}}\in U:-\varepsilon u_{\widehat{v}}^{\prime\prime}+\beta u_{\widehat{v}}^{\prime}=0,\quad u_{\widehat{v}}(-1)=\widehat{v}(1),\quad u_{\widehat{v}}(1)=\widehat{v}(2),$
implying that $\|(v,\widehat{v})\|_{V\times\mathbb{R}^{2}}$ must be equal to zero. Finally, we can prove that the inf-sup condition  (3) is satisfied with $\alpha=1$ by first noticing that it is trivially satisfied for all $u\in V$. In addition, for a given $u$ such that $\widehat{u}\neq 0$, given $v=u-u_{d}$ with $u_{d}$ any lifting of the boundary data $\widehat{u}$, and $\widehat{v}=\widehat{u}-\eta\dfrac{\widehat{u}}{\|\widehat{u}\|^{2}_{E}}$, with $\eta=-(\varepsilon u^{\prime},u_{d}^{\prime})_{0}+(\beta u^{\prime},u-u_{d})_{0}$, it holds
$\frac{a(u,(v,\widehat{v}))}{\|(v,\widehat{v})\|_{V\times\mathbb{R}^{2}}}=\|u\|_{U}.$
The result follows from noticing that the supreme over $V\times\mathbb{R}^{2}$ is always larger or equal to the left-hand side of the above equality.
With this alternative definition, it can be proved that the corresponding loss functional is given by
$\mathcal{L}^{\phi}_{r}(u_{\theta})=\|\phi\|_{V}^{2}+|u_{\theta}(-1)|^{2}+|u_{\theta}(1)|^{2}+C(u_{\theta}),$ (51)
with $\phi$ being the residual representative in $V_{M}\subseteq V$ with residual $r(u,\cdot)$ defined as in the variational formulation (42). Indeed, choosing the canonical basis for $\mathbb{R}^{2}$, the corresponding (block-diagonal) Gram matrix $\widehat{G}$ associated to the inner product inducing the norm for the space $V\times\mathbb{R}^{2}$, and the rhs ^$\widehat{\cal R}(\theta)$, respectively, have the following structure
$\widehat{G}=\left(\begin{matrix}G&\\ &I_{2}\end{matrix}\right)\quad\text{and}\quad\widehat{\cal R}(\theta)=\left(\begin{matrix}{\cal R}(\theta)\\ -u_{\theta}(-1)\\ -u_{\theta}(1)\end{matrix}\right),$ (52)
with $I_{2}$ denoting the $2\times 2$ identity matrix, $G$ the Gram matrix associated with the inner product of $V$, and ${\cal R}(\theta)$ the rhs associated with the residual $r(u_{\theta},\cdot):=\langle f\,,\cdot\rangle-(\varepsilon u^{\prime}\,,\cdot^{\prime})_{0}-(\beta u^{\prime}\,,\cdot)_{0}$ in $V^{\prime}$.
Remark 12 (Penalized weak BCs imposition).
We notice that it is also possible to consider a penalization for the weak imposition by multiplying the term $\widehat{u}\times\widehat{v}$ with a (possibly boundary-dependent) positive penalization term. This could serve as a further stabilization strategy in complex scenarios. For instance, in the advection-dominated regime.
@@ 5.3 Discrete setting
For the sake of simplicity, in all the numerical examples we assume that there is no data available for interpolation; thus, we consider $C(u_{\theta})=0$. For one-dimensional problems, we set $\Omega:=(-1,1)$ and define two kinds of test spaces. First, we consider $V_{M}$ as FE space of the standard globally continuous and piece-wise linear functions that belong to $H_{0}^{1}(\Omega)$, defined over a uniform mesh partition of $\Omega$ into $M+1$ elements $[x_{i},x_{i+1}]$, with $x_{0}=-1<x_{1}<\dots<x_{M}<x_{M+1}=1$. Similarly, we consider a 2D case where $\Omega:=(0,1)^{2}$. Here, we employ tensor-product piece-wise linear FEM test functions. The corresponding Gram matrix in both cases is constructed using the inner product, inducing the norm $\|\cdot\|_{V}$, defined in (44). Herein, we compute $G^{-1}$ explicitly in both 1D and 2D examples employing FEM test spaces.
The second test space is intended to show examples involving an orthonormal basis for the discrete space $V_{M}$. Here, we normalize with respect to the inner product (44) the first $M$ eigenfunctions of the Laplacian satisfying the Dirichlet homogeneous boundary conditions. The $m$-th eigenfunction and the corresponding orthonormal set, respectively, are given by
@@
Therefore, the loss functional (26) is explicitly given as:
@@
$\mathcal{L}^{\phi}_{r}(u_{\theta})=\dfrac{1}{\varepsilon\pi^{2}}\sum_{n=1}^{N}\sum_{m=1}^{M}\dfrac{1}{n^{2}+m^{2}}\,r\left(u_{\theta}\,,s_{m,n}\right)^{2}.$ (56)
For the architecture of the DNN approximator $u_{\theta}$, we adopt the following structure for the 1D problems: a fixed feed-forward fully connected network comprising five layers, each with 25 neurons. Throughout all layers, the activation function employed is $\tanh$. When strongly enforcing the homogeneous Dirichlet boundary conditions, we multiply the last layer’s output by $(x+1)(x-1)$. Another version includes the weak imposition of homogeneous Dirichlet boundary conditions by adding to the functional loss the following quantity (see Section 5.2):
$|u_{\theta}(-1)|^{2}+|u_{\theta}(1)|^{2}.$ (57)
As our choice for the nonlinear solver, we opt for the ADAM optimizer (refer to [35]) and initialize it with a learning rate of $0.0005$. Our nonlinear solver undergoes up to $6000$ iterations (epochs). To approximate the integral terms appearing in the loss functional definitions of the 1D problems, we employ (a) a fixed Gaussian quadrature rule with five nodes per element when considering FE functions, and (b) a trapezoidal rule with $4000$ equally-distributed nodes when considering spectral functions. Lastly, the error $|u-u_{\theta}|_{U}$ is numerically estimated using a trapezoidal rule, with $10000$ equally-distributed nodes. For the 2D example, we consider a fixed feed-forward fully connected network comprising four layers, each with 40 neurons and the same type of activation function as in the 1D cases. For the training in the spectral case, $500\times 500$ equally-distributed integration points, while equally-distributed $1000\times 1000$ integration points for estimating the error $|u-u_{\theta}|_{U}$. We strongly impose the Dirichlet boundary conditions by multiplying the last layer’s output by $x(x-1)y(y-1)$.
@@ 5.4 Smooth diffusion problems
Figure 1 shows the numerical results obtained with RVPINNs with $50$ sinusoidal test functions and strong imposition of the BCs. We observe a good performance of the RVPINNs strategy, obtaining an accurate approximation of the analytical solution, as shown in Figure 0(a). Moreover, we observe a perfect match between the quantities $\|u-u_{\theta}\|_{U}$, $\|\phi\|_{V}$, and $\sqrt{\mathcal{L}^{\phi}_{r}(u_{\theta})}$ as shown in figures 0(b) and  0(c) where, in particular, the last one shows a perfect correlation between the true error and the square root of the loss functional. Here, we compute $\|\phi\|_{V}$ employing the right-hand-side of (22). Note that $\|\phi\|_{V}$ and $\sqrt{\mathcal{L}^{\phi}_{r}(u_{\theta})}$ coincide when strongly imposing BCs. Figure 2 shows an almost identical behavior when considering $100$ FE test functions and strong imposition of BCs. We obtain similar results when considering the weak imposition of BCs, but we omit them here for brevity.
Note that in the case of pure diffusive problems, the continuity constant $\mu$ and the inf-sup stability constant $\alpha$ are equal to one when considering the $H^{1}$-seminorm (cf. (45) and (48)). Therefore, $\|\phi\|_{V}$ always defines a lower bound for the error (cf. (34)) and an upper bound up to a constant close to one if the oscillation term in (35) is sufficiently small. These theoretical considerations are confirmed in figures 1 and 2.
@@
To conclude this section, we show the results obtained for a 2D pure diffusion example, set in the domain $\Omega=(0,1)^{2}$ analogously so the analytical solution coincides with
@@
Figure 4 displays the results when we consider the spectral loss functional (55) during training with $M=N=20$ and strong imposition of the BCs. Similarly, Figure 5 displays the results when considering a $100\times 100$-element mesh with piece-wise linear FEM test functions. We observe a similar behavior, up to some oscillation due to the inexact integration, to the 1D case, showing in particular that the strategy is dimension-independent, as expected.
@@
(a) $u-\text{(best) }u_{\theta}$
(b) Estimates vs Iterations
(c) Error-Loss comparison
Figure 5: RVPINNs approximation of the 2D smooth problem; strong BCs imposition, $100\times 100$ elements mesh with piece-wise linear FEM test functions.
@@ 5.5 Delta source problem
which is only a $C^{0}(\Omega)$ solution, while the activation function $\tanh$ is smooth. Therefore, a major dominance of the oscillation term is expected, even when considering a sufficiently large number of test functions, as appreciated in Figure 6 where we have considered $100$ FE test functions and strong imposition of the BCs. Figure 7 shows similar results for the same configuration but with the weak imposition of the BCs. In both cases, a good representation of the analytical solution is obtained despite being approximated with a smooth function. Additionally, figures 6 and  7 show that the loss functional defines a robust estimation of the error in the pre-asymptotic regime (before the dominance of the oscillation term). A similar behavior is obtained when considering the loss functional with spectral test functions– omitted here for brevity.
@@
Figure 7: RVPINNs approximation of the delta diffusion problem; weak BCs imposition, $100$ FE test functions.
@@ 5.6 Advection-dominated diffusion problem
exhibiting a strong gradient near the right boundary when $\varepsilon\rightarrow 0$ (cf. figures 7(a) and 9(a)). We restrict the discussion to spectral test functions (figures  8– 10), as we observe similar behaviors when considering FE test functions.
Figure 8 shows the results obtained for $\varepsilon=0.1$, considering $50$ test functions and strong imposition of the BCs. Contrary to the pure diffusive case, Figure 7(b) reflects that the lines representing the error and the square root of the loss functional (respectively, the norm of the discrete Riesz representative with respect to the norm of the test space $V\times\mathbb{R}^{2}$) do not overlap. This behavior is expected as, in this case, the continuity constant $\mu$ is not equal to one. Indeed, since the Poincaré constant is well-known in 1D, we can explicitly calculate it: $\mu=1+2/(\varepsilon\pi)$. It deteriorates when $\varepsilon\rightarrow 0$, i.e., the estimations for this particular setting are not robust in terms of the physical parameter $\varepsilon$. Nevertheless, as we can see in Figure 7(c), our theoretical a posteriori error bounds  (34) and (35) are satisfied. We also observe a decay in the expected optimal correlation between the error and the square root of the loss starting near iteration $3000$, apparently due to stability issues, that is in agreement with the oscilation term appearing in the a posteriori error estimate (35). This suboptimality is improved when considering the weak imposition of the boundary conditions, as shown in Figure 9. This weak imposition of BCs could thus serve as a stabilization technique. This is further confirmed in Figure 10, where we show the results obtained for $\varepsilon=0.005$, with $200$ test basis functions and weak BCs. When considering instead the strong imposition of the BCs and the same number of test functions, we do not observe convergence within the first $6000$ iterations–omitted here for brevity. We observe that when considering the weak BCs, the loss functional exhibits some differences with the norm of the Riesz representative in $V$, as we intentionally plot the incomplete estimation to highlight the contribution of weak BC imposition in $\mathbb{R}^{2}$ of the full Riesz representative (cf. Eq. (52)). Nevertheless, in this case, we still observe a good correlation between the loss and the true error, as expected.
To guarantee good estimations in RVPINNs, we need adequate variational formulations and norms for the trial and test spaces, so the resulting inf-sup and boundedness constants are independent of the physical parameters. To exemplify this with a numerical example, consider the following (weighted) alternative variational formulation (with $\Omega:=(-1,1)$)
Find $\text{Find }u\in U:=H_{0}^{1}(\Omega):(\varepsilon\,e^{-1/\varepsilon}u^{\prime}\,,\,v^{\prime})_{0}=(e^{-1/\varepsilon}f\,,\,v)_{0},\quad\forall\,v\in V:=U,$
where the last can be deduced from (41), by first multiplying by $e^{-1/\varepsilon}v$ (assuming enough regularity for the solution $u$), and then performing integration by parts for the term involving the diffusion term. This alternative variational formulation is bounded and coercive, with constants $\mu=\alpha=1$, with respect to the norms:
$\|\cdot\|_{U}^{2}:=\|\cdot\|_{V}^{2}:=(\varepsilon\,e^{-1/\varepsilon}\cdot^{\prime}\,,\,\cdot^{\prime})_{0},$
expecting a perfect correlation of the involved quantities as in the pure diffusive case. Indeed, Figure 11 shows the numerical results obtained for $\varepsilon=0.5$, considering 200 piece-wise linear functions, confirming the robustness of RVPINNs.
@@
Figure 9: RVPINNs approximation of the diffusion-advection problem with $\varepsilon=0.1$; weak BCs imposition, $50$ spectral test functions.
@@
Figure 10: RVPINNs approximation of the diffusion-advection problem with $\varepsilon=0.005$; weak BCs imposition, $200$ spectral test functions.
(a) DNN best approximation
(b) Estimates vs Iterations
(c) Error-Loss comparison
Figure 11: RVPINNs approximation of the diffusion-advection problem with $\varepsilon=0.5$; weighted variational formulation, strong BCs imposition, $200$ linear FEM test functions.